# On determinant of checkerboard colorable virtual knots

### v1

@@ Abstract.
For classical knots, it is well known that their determinants mod $8$ are classified by the Arf invariant. Boden and Karimi introduced a determinant of checkerboard colorable virtual knots. We prove that their determinant mod $8$ is classified by the coefficient of $z^{2}$ in the ascending polynomial which is an extension of the Conway polynomial for classical knots.
@@ 1. Introduction
For a knot in $S^{3}$, let $\Delta_{K}(t)$ denote the Conway-normalized Alexander polynomial, that is, $\Delta_{K}(t)=\det(t^{1/2}V-t^{-1/2}V^{T})$ for a Seifert matrix $V$ of $K$. The absolute value $\det K=|\Delta_{K}(-1)|$ is called the determinant of $K$. Here $\Delta_{K}(t)$ is related to the Conway polynomial $\nabla_{K}(z)$ of $K$ via $\Delta_{K}(t)=\nabla_{K}(t^{-1/2}-t^{1/2})$. Let $\operatorname{Arf}K\in\mathbb{Z}/2\mathbb{Z}$ denote the Arf invariant of $K$. Then, it is well known that
@@
Figure 1. A virtual knot diagram, the corresponding Gauss diagram, and (projection of) knot in a thickened torus.
@@
Boden, Chrisman, and Karimi [1] introduced a determinant of a pair $(L,F)$, where $L$ is a link in $\Sigma\times[0,1]$ with $[L]=0$ in $H_{1}(\Sigma\times[0,1];\mathbb{Z}/2\mathbb{Z})$ and $F$ is a spanning surface for $L$. Their determinant can be applied to checkerboard colorable virtual links introduced by Kamada [13] since they admit spanning surfaces. Boden and Karimi [4] defined another determinant for these links, which is independent of the choice of a spanning surface. Now, it is natural to expect that the Arf invariant for classical knots can be extended to checkerboard colorable virtual knots so that the extension and the determinant in [4] fit into (1.1).
When a knot $K$ satisfies $[K]=0$ in $\mathbb{Z}$-coefficients and $F$ is a Seifert surface of $K$, the Arf invariant is defined for $(K,F)$ and used as an obstruction for virtual concordance of Seifert surfaces by Chrisman, Mukherjee [10] and Boden, Karimi [3]. The Arf invariant for $(K,F)$, however, does not satisfy (1.1) above (see Example 3.5). We here focus on the fact that $\operatorname{Arf}K=v_{2}(K)\bmod 2$ for classical knots $K$, where $v_{2}(K)$ denotes the coefficient of $z^{2}$ in the Conway polynomial $\nabla_{K}(z)$. For a Gauss diagram $G$ of a virtual knot with a basepoint $b$ on the circle disjoint from chords, let $\nabla_{\mathrm{asc}}(G_{b})(z)$ (resp. $\nabla_{\mathrm{des}}(G_{b})(z)$) denote the ascending (resp. descending) polynomial introduced by Chmutov, Khoury, and Rossi [7]. For classical knots, $\nabla_{\mathrm{asc}}(G_{b})(z)$ and $\nabla_{\mathrm{des}}(G_{b})(z)$ do not depend on the choice of a diagram and basepoint, and they coincide with the Conway polynomial.
We write $v_{2,1}(G_{b})$ (resp. $v_{2,2}(G_{b})$) for the coefficient of $z^{2}$ in $\nabla_{\mathrm{asc}}(G_{b})(z)$ (resp. $\nabla_{\mathrm{des}}(G_{b})(z)$) and show that they behave well for a certain class of virtual knots containing checkerboard colorable virtual knots as the case $p=2$.
@@
Let $G$ be a Gauss diagram of a mod $p$ almost classical knot. Then, for $j=1,2$, $v_{2,j}(G_{b})\bmod p$ do not depend on the choice of a basepoint $b$. Moreover, $v_{2,1}(G_{b})\equiv v_{2,2}(G_{b})\bmod p$.
@@
Note here that a similar result for almost classical knots is mentioned in [3, Section 4], but the definition of the determinant is different from [4] and ours. This difference arises from that of Gordon-Litherland linking forms (see a discussion after [5, Definition 3.1]).
By Corollary 1.2, while the Arf invariant derived from Seifert matrices coincides with the coefficient of $z^{2}$ in the Conway polynomial mod $2$ in classical knot theory, they play different roles in virtual knot theory. The former is an obstruction for concordance and the latter controls the determinant mod $8$. Finally, this paper is based on the first author’s master thesis at Hiroshima University written in Japanese.
@@ Acknowledgments
The authors would like to thank Noboru Ito, Kodai Wada, and Yuka Kotorii for their helpful comments. This study was supported in part by JSPS KAKENHI Grant Numbers JP20K14317 and JP23K12974.
@@ 2. Preliminaries
Moreover, an oriented virtual link $L$ is said to be mod $p$ almost classical if $L$ admits a mod $p$ Alexander numberable diagram. As special cases, a mod $2$ (resp. mod 00) almost classical link is also said to be checkerboard colorable in [13] (resp. almost classical in [18]).
@@
Note that a checkerboard colorable virtual link actually admits a checkerboard coloring in the sense of [13, Section 2] or [8, Introduction].
Example 2.2.
Recall here that every oriented virtual link is represented by a Gauss diagram, which is a chord diagram with oriented chords (called arrows) and with signs $+1$ or $-1$ assigned to each chord. See, for example, [7]. Throughout this paper, orientations of circles of Gauss diagrams (and arrow diagrams in Section 3) are assumed to be counterclockwise.
We recall from [2, Section 1] the definition of an index of a chord.
Definition 2.3.
Let $G$ be a Gauss diagram of a virtual knot. The index $I(c)\in\mathbb{Z}$ of a chord $c$ of $G$ with sign $\varepsilon$ is defined by $I(c)=\varepsilon(r_{+}-r_{-}+l_{-}-l_{+})$, where $r_{\pm}$ (resp. $l_{\pm}$) denote the numbers of chords with sign $\pm$ intersecting $c_{i}$ from left to right (resp. from right to left) when the head of $c$ is oriented up.
The following lemma is stated after Definition 5.1 in [2].
Lemma 2.4.
Let $G$ be a Gauss diagram of oriented virtual knot diagram $D$. Then $D$ is mod $p$ Alexander numberable if and only of $I(c)\equiv 0\bmod p$ for any chord $c$ of $G$.
The next result is shown in much the same way as [2, Theorem 6.1]. See also an argument written before Theorem 5.2 in [2].
@@
    $[L]\in H(\Sigma\times[0,1];\mathbb{Z}/p\mathbb{Z})$ is trivial, where $\Sigma$ is a Carter surface of $K$.
@@ 2.2. Determinant for checkerboard colorable links
We recall from [4, Section 3] the definition of the coloring matrix.
@@
$b_{ij}=\begin{cases}2&\text{if $a_{j}$ is the over-crossing arc at $c_{i}$,}\\ -1&\text{if $a_{j}$ is one of an under-crossing arc at $c_{i}$,}\\ 0&\text{otherwise.}\end{cases}$
@@
Therefore, $\det K=1$ (see Definition 2.8).
@@
The determinant $\det L$ of checkerboard colorable virtual link $L$ is defined to be the absolute value of the determinant of an $(n-1)\times(n-1)$ minor of $B(D)$, which is well-defined due to [4, Proposition 3.1].
To show a kind of skein relation for $\det L$ in the next subsection, we give another interpretation of $\det L$ in terms of a mock Seifert matrix introduced in [5, Definition 3.5]. Let $\Sigma$ be an orientable closed surface. For disjoint oriented virtual knots $K_{1}$ and $K_{2}$ in $\Sigma\times[0,1]$, we define the virtual linking number $\mathrm{vlk}(K_{1},K_{2})\in\mathbb{Z}$ by the identity
$[K_{2}]=\mathrm{vlk}(K_{1},K_{2})[\mu]\in H_{1}(\Sigma\times[0,1]\setminus K_{1},\Sigma\times\{1\};\mathbb{Z})\cong\mathbb{Z}.$
We recall the Gordon-Litherland form $\mathcal{L}_{F}$ from [5]. Let $\widetilde{F}$ be the boundary of a neighborhood of $F\subset\Sigma\times[0,1]$ and $\pi\colon\widetilde{F}\to F$ be the induced double cover. One can define a transfer map $\tau\colon H_{1}(F;\mathbb{Z})\to H_{1}(\widetilde{F};\mathbb{Z})$, and then the Gordon-Litherland form $\mathcal{L}_{F}\colon H_{1}(F)\times H_{1}(F)\to\mathbb{Z}$ is defined by $\mathcal{L}_{F}(\alpha,\beta)=\mathrm{vlk}(\tau(\alpha),\beta)$.
Definition 2.9.
Let $L$ be a link in $\Sigma\times[0,1]$ with $[L]=0\in H_{1}(\Sigma\times[0,1];\mathbb{Z}/2\mathbb{Z})$ and let $F$ be a spanning surface for $L$, that is, $F$ is (not necessarily orientable) connected compact surface whose boundary is $L$. A matrix for the Gordon-Litherland form $\mathcal{L}_{F}$ associated with $F$ is called a mock Seifert matrix for $L$.
Let $L$ be a checkerboard colorable virtual link and $D$ a diagram of $L$. We write for $X$ the quotient space obtained from $\Sigma\times[0,1]$ by identifying $\Sigma\times\{1\}$ with a point. Let $X_{2}$ be the double cover of $X$ branched over $L$ and $X_{\infty}$ the infinite cyclic cover of $X\setminus L$. It is shown in [4, Section 3] that we can compute the first elementary ideal $\mathcal{E}_{1}$ of the Alexander module $H_{1}(X_{\infty};\mathbb{Z})$ over the ring $\mathbb{Z}[t,t^{-1}]$ from a matrix $A(D)$ obtained by Fox’s free derivative. As mentioned before Proposition 3.1 in [4], the coloring matrix $B(D)$ is obtained from $A(D)$ by substituting $t=-1$ (up to sign for any given row). Therefore, one can compute the first elementary ideal of $H_{1}(X_{2};\mathbb{Z})$ over the ring $\mathbb{Z}[-1]=\mathbb{Z}$ from $B(D)$. On the other hand, it is shown in [5, Theorem 3.9] that a mock Seifert matrix for $L$ is a presentation matrix of $H_{1}(X_{2};\mathbb{Z})$. Therefore, we obtain the following consequence (see [5, Remark 3.11]).
@@
For almost classical oriented link $L$, the Alexander polynomial $\Delta_{L}(t)$ is defined in [2, Section 7] (see also [15]). It is shown in [4, Proposition 4.1] that $\det L=|\Delta_{L}(-1)|$ holds.
@@
Let $K$ be the virtual knot $6.87548$ in Green’s table [12] (see also [2, Figure 20]). Using a Seifert surface in [9, Figure 14] as a spanning surface, we have
@@
which is also obtained by adding $V^{+}$ and $V^{-}$ in [9, Example (6.87548)]. Then one has $\det K=1$. Note that this result is compatible with [2, Table 2].
@@ 2.3. Skein relation
We prove a kind of skein relation for $\det L$ similar to [2, Theorem 7.11] for the Alexander polynomial of almost classical links. The idea of the proof is based on [11, Lemma 3].
@@
Let $L_{+}$, $L_{-}$ and $L_{0}$ be oriented virtual links in Figure 4. Suppose that one of them is checkerboard colorable (then the others are also checkerboard colorable) and consider spanning surfaces in Figure 5. Then, mock Seifert matrices $S_{+}$, $S_{-}$, $S_{0}$ associated with the spanning surfaces satisfy
@@
Figure 4. Oriented virtual links $L_{+}$, $L_{-}$, $L_{0}$ that are identical except within a $3$-ball.
Figure 5. Spanning surfaces $F_{+}$, $F_{-}$, $F_{0}$ that are identical except within a $3$-ball.
@@
Fix a basis of $H_{1}(F_{0};\mathbb{Z})$ and let $S_{0}$ be the corresponding mock Seifert surface for $F_{0}$. Since $F_{+}$ is obtained from the connected surface $F_{0}$ by attaching a band, there exists a loop $\ell$ on $F_{+}$ passing through the band. By the Mayer-Vietoris exact sequence, the union of the basis of $H_{1}(F_{0};\mathbb{Z})$ and $[\ell]\in H_{1}(F_{+};\mathbb{Z})$ is a basis of $H_{1}(F_{+};\mathbb{Z})$. Now, the union also gives a basis of $H_{1}(F_{-};\mathbb{Z})$. Let $S_{\pm}$ be mock Seifert surfaces for $F_{\pm}$ with respect to the bases. Then one can see that

### v2

# On the determinant of checkerboard colorable virtual knots

@@ Abstract.
For classical knots, Murasugi showed that the determinant modulo $8$ is classified by the Arf invariant. Boden and Karimi introduced a determinant for checkerboard colorable virtual knots. We prove that this determinant modulo $8$ is classified by the coefficient of $z^{2}$ in the ascending polynomial, an extension of the Conway polynomial for classical knots.
@@ 1. Introduction
For a knot in $S^{3}$, let $\Delta_{K}(t)$ denote the Conway-normalized Alexander polynomial, that is, $\Delta_{K}(t)=\det(t^{1/2}V-t^{-1/2}V^{T})$ for a Seifert matrix $V$ of $K$. The special value $\Delta_{K}(-1)$ is known to be an odd integer, and its absolute value $\det K$ is called the determinant of $K$. Here $\Delta_{K}(t)$ is related to the Conway polynomial $\nabla_{K}(z)$ of $K$ via $\Delta_{K}(t)=\nabla_{K}(t^{-1/2}-t^{1/2})$. Let $\operatorname{Arf}K\in\mathbb{Z}/2\mathbb{Z}$ denote the Arf invariant of $K$. Murasugi [15] established the following relation, which is now well known:
@@
Figure 1. A virtual knot diagram, the corresponding Gauss diagram, and (projection of) a knot in a thickened torus.
@@
Boden, Chrisman, and Karimi [1] introduced a determinant of a pair $(L,F)$, where $L$ is a link in $\Sigma\times[0,1]$ with $[L]=0$ in $H_{1}(\Sigma\times[0,1];\mathbb{Z}/2\mathbb{Z})$ and $F$ is a spanning surface for $L$. Their determinant can be applied to checkerboard colorable virtual links introduced by Kamada [13] since they admit spanning surfaces. Boden and Karimi [3] defined another determinant for these links, which is independent of the choice of a spanning surface. Now, it is natural to expect that the Arf invariant for classical knots can be extended to checkerboard colorable virtual knots so that the extension and the determinant in [3] fit into the equality (1.1).
When a knot $K$ satisfies $[K]=0$ in $\mathbb{Z}$-coefficients and $F$ is a Seifert surface of $K$, the Arf invariant is defined for $(K,F)$ and used as an obstruction to virtual concordance of Seifert surfaces by Chrisman, Mukherjee [10] and Boden, Karimi [5]. The Arf invariant for $(K,F)$, however, does not satisfy (1.1) above (see Example 3.5).
We here focus on the fact that $\operatorname{Arf}K=v_{2}(K)\bmod 2$ for classical knots $K$, where $v_{2}(K)$ denotes the coefficient of $z^{2}$ in the Conway polynomial $\nabla_{K}(z)$. For a Gauss diagram $G$ of a virtual knot with a basepoint $b$ on the circle disjoint from chords, let $\nabla_{\mathrm{asc}}(G_{b})(z)\in\mathbb{Z}[z]$ (resp. $\nabla_{\mathrm{des}}(G_{b})(z)$) denote the ascending (resp. descending) polynomial introduced by Chmutov, Khoury, and Rossi [7]. For classical knots, $\nabla_{\mathrm{asc}}(G_{b})(z)$ and $\nabla_{\mathrm{des}}(G_{b})(z)$ do not depend on the choice of a Gauss diagram and basepoint, and they coincide with the Conway polynomial.
We write $v_{2,1}(G_{b})$ (resp. $v_{2,2}(G_{b})$) for the coefficient of $z^{2}$ in $\nabla_{\mathrm{asc}}(G_{b})(z)$ (resp. $\nabla_{\mathrm{des}}(G_{b})(z)$) and show that they behave well for a certain class of virtual knots containing checkerboard colorable virtual knots.
@@
Let $p$ be a non-negative integer and let $G$ be a Gauss diagram of a mod $p$ almost classical knot. Then, for $j=1,2$, $v_{2,j}(G_{b})\bmod p$ do not depend on the choice of a basepoint $b$. Moreover, $v_{2,1}(G_{b})\equiv v_{2,2}(G_{b})\bmod p$.
@@
Note here that a similar result for almost classical knots is mentioned in [5, Section 4], but the definition of the determinant is different from [3] and ours. This difference arises from the difference in Gordon-Litherland linking forms (see a discussion after [4, Definition 3.1]).
By Corollary 1.2, while the Arf invariant derived from Seifert matrices coincides with the coefficient of $z^{2}$ in the Conway polynomial modulo $2$ in classical knot theory, they play different roles in virtual knot theory. The former is an obstruction to concordance, and the latter controls the determinant modulo $8$. Finally, this paper is based on the first author’s master thesis at Hiroshima University, written in Japanese.
@@ Acknowledgments
The authors are grateful to Kodai Wada for his careful reading of the manuscript, and to Noboru Ito and Yuka Kotorii for their helpful comments. They also wish to express their thanks to the anonymous referee for valuable comments. This study was supported in part by JSPS KAKENHI Grant Numbers JP20K14317 and JP23K12974.
@@ 2. Preliminaries
Moreover, an oriented virtual link $L$ is said to be mod $p$ almost classical if $L$ admits a mod $p$ Alexander numberable diagram. As special cases, a mod $2$ almost classical link is said to be checkerboard colorable in [13], and mod 0 almost classical link is said to be almost classical in [20].
@@
Being mod $p$ almost classical is characterized by the first homology group of a thickened surface as stated in the next theorem, which is shown in much the same way as [2, Theorem 6.1]. See also an argument written before Theorem 5.2 in [2].
@@
    $L]\in H(\Sigma\times[0,1];\mathbb{Z}/p\mathbb{Z})$ is trivial, where $\Sigma$ is the Carter surface of a diagram of $L$.
In particular, a checkerboard colorable virtual link $L$ actually admits a checkerboard coloring in the sense of [13, Section 2] (or [8, Introduction]), and it has a spanning surface in $\Sigma\times[0,1]$.
Example 2.3.
Recall here that every oriented virtual link is represented by a Gauss diagram, which is a chord diagram with oriented chords (called arrows) and with signs +1+1 or 1-1 assigned to each chord. See, for example, [7]. Throughout this paper, orientations of circles of Gauss diagrams (and arrow diagrams in Section 3) are assumed to be counterclockwise.
We recall from [2, Section 1] the definition of the index of a chord. See also [18].
Definition 2.4.
Let $G$ be a Gauss diagram of a virtual knot. The index $I(c)\in\mathbb{Z}$ of a chord $c$ of $G$ with sign $\varepsilon$ is defined by $I(c)=\varepsilon(r_{+}-r_{-}+l_{-}-l_{+})$, where $r_{\pm}$ (resp. $l_{\pm}$) denote the numbers of chords with sign $\pm$ intersecting $c$ from left to right (resp. from right to left) when the head of $c$ is oriented up.
The following lemma is stated after Definition 5.1 in [2].
Lemma 2.5.
Let $G$ be a Gauss diagram of an oriented virtual knot diagram $D$. Then $D$ is mod $p$ Alexander numberable if and only of $I(c)\equiv 0\bmod p$ for any chord $c$ of $G$.
@@ 2.2. Determinant for checkerboard colorable links
We recall from [3, Section 3] the definition of the coloring matrix.
@@
[…]
Definition 2.7.
The determinant $\det L$ of checkerboard colorable virtual link $L$ is defined to be the absolute value of the determinant of an $(n-1)\times(n-1)$ minor of $B(D)$, which is well-defined due to [3, Proposition 3.1].
@@
Therefore, $\det K=1$.
To show a kind of skein relation for $\det L$ in the next subsection, we give another interpretation of $\det L$ in terms of a mock Seifert matrix introduced in [4, Definition 3.5]. Let $\Sigma$ be an orientable closed surface. For two disjoint oriented knots $K_{1}$ and $K_{2}$ in $\Sigma\times[0,1]$, we define the virtual linking number $\mathrm{vlk}(K_{1},K_{2})\in\mathbb{Z}$ by the identity
$K_{2}]=\mathrm{vlk}(K_{1},K_{2})[\mu]\in H_{1}(\Sigma\times[0,1]\setminus K_{1},\Sigma\times\{1\};\mathbb{Z})\cong\mathbb{Z}.$
[…]
@@
Let $L$ be a link in $\Sigma\times[0,1]$ with $L]=0\in H_{1}(\Sigma\times[0,1];\mathbb{Z}/2\mathbb{Z})$ and let $F$ be a spanning surface for $L$, that is, $F$ is (not necessarily orientable) connected compact surface whose boundary is $L$. A matrix for the Gordon-Litherland form $\mathcal{L}_{F}$ associated with $F$ is called a mock Seifert matrix for $L$.
Let $L$ be a checkerboard colorable virtual link, $D$ a diagram of $L$, and $\Sigma$ the Carter surface of $D$. We write for $X$ the quotient space obtained from $\Sigma\times[0,1]$ by identifying $\Sigma\times\{1\}$ with a point. Let $X_{2}$ be the double cover of $X$ branched over $L$ and $X_{\infty}$ the infinite cyclic cover of $X\setminus L$. It is shown in [3, Section 3] that we can compute the first elementary ideal $\mathcal{E}_{1}$ of the Alexander module $H_{1}(X_{\infty};\mathbb{Z})$ over the ring $\mathbb{Z}[t,t^{-1}]$ from a matrix $A(D)$ obtained by Fox’s free derivative. As mentioned before Proposition 3.1 in [3], the coloring matrix $B(D)$ is obtained from $A(D)$ by substituting t=1t=-1 (up to sign for any given row). Therefore, one can compute the first elementary ideal of $H_{1}(X_{2};\mathbb{Z})$ over the ring $\mathbb{Z}[-1]=\mathbb{Z}$ from $B(D)$. On the other hand, it is shown in [4, Theorem 3.9] that a mock Seifert matrix for $L$ is a presentation matrix of $H_{1}(X_{2};\mathbb{Z})$. Therefore, we obtain the following consequence (see [4, Remark 3.11]).
@@
For an almost classical oriented link $L$, the Alexander polynomial $\Delta_{L}(t)$ is defined in [2, Section 7] (see also [16]). It is shown in [3, Proposition 4.1] that $\det L=|\Delta_{L}(-1)|$ holds.
Figure 4. The virtual knot $6.87548$ in Green’s table [12] (see [2, Figure 20] for its Gauss diagram).
@@
Let $K$ be the virtual knot illustrated in Figure 4. Using a Seifert surface in [9, Figure 14] as a spanning surface, we have
@@
which is also obtained by adding $V^{+}$ and $V^{-}$ in [9, Example (6.87548) in Section 6]. Then one has $\det K=1$. Note that this result is compatible with [2, Table 2].
@@ 2.3. Skein relation
We prove a kind of skein relation for $\det L$ similar to [2, Theorem 7.11] for the Alexander polynomial of almost classical links. The idea of the proof is based on [11, Lemma 3]. Recall from Theorem 2.2 that a checkerboard colorable virtual link has a spanning surface in $\Sigma\times[0,1]$.
@@
Let $_{+}, LL_{-}$ and $L_{0}$ be oriented virtual links in Figure 5. Suppose that one of them is checkerboard colorable (then the others are also checkerboard colorable) and consider spanning surfaces in Figure 6. Then, mock Seifert matrices $_{+}, SS_{-}$, $S_{0}$ associated with the spanning surfaces satisfy
@@
Figure 5. Oriented virtual links $_{+}, LL_{-}$, $L_{0}$ that are identical except within a 33-ball.
Figure 6. Spanning surfaces $_{+}, FF_{-}$, $F_{0}$ that are identical except within a 33-ball.
@@
Fix a basis of $H_{1}(F_{0};\mathbb{Z})$ and let $S_{0}$ be the corresponding mock Seifert surface for $F_{0}$. Since $F_{+}$ is obtained from the connected surface $F_{0}$ by attaching a band, there exists a loop $\ell$ on $F_{+}$ passing through the band. By the Mayer-Vietoris exact sequence, the union of the basis of $H_{1}(F_{0};\mathbb{Z})$ and $\ell]\in H_{1}(F_{+};\mathbb{Z})$ is a basis of $H_{1}(F_{+};\mathbb{Z})$. Now, the union also gives a basis of $H_{1}(F_{-};\mathbb{Z})$. Let $S_{\pm}$ be mock Seifert surfaces for $F_{\pm}$ with respect to the bases. Then one can see that